OpenStudy (anonymous):

The monthly sales (in thousands of units) of a seasonal product are approximated by S= 74.50 + 43.75 sin pi t/6 Where t is time in months, with t=1 corresponding to January. Determine the months when sales exceed 100,000 units

OpenStudy (anonymous):

they really reach for these word problems don't they?

OpenStudy (anonymous):

Lol yeah. It's terrible.

OpenStudy (anonymous):

\[74.50+43.75\sin(\frac{\pi t}{6})>100\] \[43.75\sin(\frac{\pi t}{6})>25.5\] \[\sin(\frac{\pi t}{6})>.583\] rounded. we can try various values for t since they all have too be integers

OpenStudy (anonymous):

okay thanks!

OpenStudy (anonymous):

1 doesn't work, because \(\sin(\frac{\pi}{6})=.5\) but 2, 3 4 will work

OpenStudy (anonymous):

can visualize it here http://www.wolframalpha.com/input/?i=sin%28pi+t%2F6%29%3E.583+++from+1+to+12